2(y-x)=5+2x
2(y+x)=5+2y
Sytems..!!

Answer by checkley77(12844)

2(y-x)=5+2x
--------------------------------
2(y+x)=5+2y
2y+2x=5+2y
2y-2y=-2x+5
0=-2x+5
2x=5
x=5/2
x=2.5 ans.
2(y-(2.5))=5+2(2.5)
2(y-2.5)=5+5
2y-5=10
2y=10+5
y=15/2
y=7.5 ans.
PROOF:
2(7-2.5)=5+2*2.5
2*5=5+5
10=10